OpenStudy (anonymous):

Factor completely 3x2 + 9x – 54 (3x – 9)(x + 6) 3(x – 3)(x + 6) (3x – 6)(x + 9) 3(x – 2)(x + 9)

OpenStudy (anonymous):

4th one

OpenStudy (anonymous):

= 3(x^2 + x - 18) = 3(x-2)(x+9) D..

OpenStudy (anonymous):

thanks guys!